OpenStudy (anonymous):

what quadratic equation has solution 3+i and 3-i

OpenStudy (mertsj):

(x-(3+i)][x-(3-i)]=0

jhonyy9 (jhonyy9):

so x2 -(3-i)x-(3+i)x+(9+1)=x2 -x(3-i+3+i)+10=x2 -6x +10